# Semiparametric Copula Estimation for Spatially Correlated Multivariate Mixed Outcomes: Analyzing Visual Sightings of Fin Whales from Line Transect Survey

### v2

@@ Introduction
To model the interactions among variables, multivariate analysis is an ubiquitous tools, with many methodologies developed, such as factor analysis, principal component analysis (e.g., Manly and Alberto, 2016), graphical modeling (e.g., Jordan, 2004), and copulas (e.g., Joe, 2014). In particular, copulas are attractive tools for explicitly modeling dependence structures without modeling marginal distributions and can also be successfully applied even when outcomes are mixed (i.e., outcomes include both continuous and discrete variables). As techniques for estimating copulas, Hoff (2007) proposed the extended rank likelihood for parametric copulas and developed a simple Gibbs sampler for Bayesian inference on copulas. Although these methods are useful for modeling the dependence structure, multivariate observations are assumed to be independent of each other. However, beginning with the example of marine ecosystems mentioned above, in many applications such as epidemiology, climatology, medicine, and sociology, multivariate data with location information are often available, and spatial correlation must be appropriately considered; otherwise, the estimation of the dependence structure among outcomes could be severely biased.
@@
Related work in spatial modeling includes several studies on modeling multivariate spatial data. For example, Dey et al. (2022) and Krock et al. (2023) proposed efficient modeling (relatively high-dimensional) of multivariate spatial data for Gaussian outcomes, whereas Feng and Dean (2012) and Torabi (2014) considered multivariate models based on generalized linear mixed models. Furthermore, attempts have been made to use copulas in multivariate spatial modeling, such as Musafer et al. (2017), Krupskii et al. (2018), Krupskii and Genton (2019), and Gong and Huser (2022). Because the above methods consider the joint estimation of both marginal and dependence structures, one may lose the efficiency of the estimation by a large number of nuisance parameters in the marginal distributions, while the estimation of the dependence could be biased owing to potential specifications of the marginal distributions. Hence, it is more reasonable to model the dependence structure directly, as in the proposed method.
@@
The remainder of this paper is organized as follows. Section 2 introduces the extended rank likelihood and semiparametric Bayesian inference for the dependence structure proposed by Hoff (2007). Section 3 proposes a semiparametric Bayesian inference for the dependence structure in spatially correlated mixed outcomes and develops an efficient algorithm for computing the posterior using the MCMC. The algorithm is then extended to a more scalable implementation using the nearest-neighbor Gaussian process, even for large spatial datasets. Section 4 presents the simulation results, including a comparison with the method of Hoff (2007), to validate the usefulness of the proposed method. Section 5 presents the application of the proposed method using real data collected during an international synoptic krill survey in the Scotia Sea of the Antarctic Peninsula. Section 6 provides directions for future research. The R code for implementing the proposed method is available in the GitHub repository (https://github.com/t-momozaki/spBGC).
Extended rank likelihood for semiparametric copula estimation
@@
Suppose that we observe $p$-dimensional observations of mixed outcomes $\bm{y}_{i}=(y_{i1},\ldots,y_{ip})^{\top}$ for $i=1,2,\ldots,n$, where each variable could be a variety of outcomes, such as continuous, count, and ordered variables. Here, we are interested in the dependence structure of $p$ outcomes. The marginal distributions of each component in $\bm{y}_{i}$ are not of interest and are the nuisance parameters in Hoff (2007) and our framework. Let $\bm{z}_{i}$ be a $p$-dimensional continuous latent variable for $\bm{y}_{i}$. Then, we consider the following Gaussian copula model:
@@
where $\bm{R}$ is the $p\times p$ correlation matrix and $\bm{I}_{n}$, $\otimes$, $F_{j}^{-1}$, and $\Phi(\cdot)$ denote the $n\times n$ identity matrix, Kronecker product, inverse of an unknown univariate cumulative distribution function, and cumulative distribution function of the standard normal distribution, respectively. Here, we focus on correlation matrix $\bm{R}$ as the dependence structure. In this model, Hoff (2007) defines the following extended rank likelihood:
[…]
where $\phi_{k}(\bm{x};\bm{\mu},\bm{\Sigma})$ is the $k$-dimensional multivariate normal density with a mean vector $\bm{\mu}$ and a variance-covariance matrix $\bm{\Sigma}$. Observing $\bm{y}=(\bm{y}_{1}^{\top},\bm{y}_{2}^{\top},\ldots,\bm{y}_{n}^{\top})^{\top}$ indicates that $\bm{z}$ must lie within the set
$D=\{\bm{z}\in\mathbb{R}^{pn}:\max(z_{kj}:y_{kj}<y_{ij})<z_{ij}<\min(z_{kj}:y_{ij}<y_{kj})\}$
as $F_{j}$ is a nondecreasing function and observing $y_{ij}<y_{i^{\prime}j}$ for any $i\neq i^{\prime}$ implies that $z_{ij}<z_{i^{\prime}j}$. Note that this likelihood depends only on the parameter of interest, $\bm{R}$, and not on the nuisance parameters, $\bm{F}=(F_{1},F_{2},\ldots,F_{p})^{\top}$. It can be regarded as a type of marginal likelihood in the presence of nuisance parameters and can be derived in terms of the decomposition theorem on sufficient statistics. That is,
$p(\bm{y}|\bm{R},\bm{F})$ $=p(\bm{z}\in D,\bm{y}|\bm{R},\bm{R})$
@@
and to estimate $\bm{R}$, it is sufficient to use only $\Pr(\bm{z}\in D|\bm{R})$. Function (2) is called the extended rank likelihood because it is derived from the marginal probability of the ranks, and can be seen as a multivariate version of the rank likelihood (Pettitt, 1982; Heller and Qin, 2001), being free of the nuisance parameters for continuous and discrete data. For a more detailed discussion on the extended rank likelihood and sufficient statistics, see Hoff (2007), Section 5.
@@
Copula estimation under spatial correlation
Latent models with spatial correlation
Suppose that we obtain observations of mixed outcomes $\bm{y}_{\mathcal{S}}$ equipped with location information $\mathcal{S}$, where $\bm{y}_{\mathcal{S}}=(\bm{y}(\bm{s}_{1})^{\top},\bm{y}(\bm{s}_{2})^{\top},\ldots,\bm{y}(\bm{s}_{n})^{\top})^{\top}$, $\bm{y}(\bm{s}_{i})=(y_{1}(\bm{s}_{i}),y_{2}(\bm{s}_{i}),\ldots,y_{p}(\bm{s}_{i}))^{\top}$, $\mathcal{S}=\{\bm{s}_{1},\bm{s}_{2},\ldots\bm{s}_{n}\}$, and $\bm{s}_{i}$ denotes a two-dimensional vector of longitude and latitude. These observations are characterized by spatial dependence and correlation, and nearby observations have similar properties. In such data, spatial correlations should be adequately considered; otherwise, the estimation of the dependence structure among outcomes would be severely biased. Therefore, we consider a hierarchical spatial model and latent multivariate Gaussian process. We define the following spatial Gaussian copula as a Gaussian copula model combined with a spatial Gaussian process to consider spatial correlation:
@@
    $\mathcal{S}_{-i}=\{\bm{s}_{i^{\prime}}|i\neq i^{\prime},i^{\prime}\in\mathbb{N}\}$, $\bm{H}_{\bm{s}_{i},\mathcal{S}_{-i}}$ is the $(n-1)$-dimensional cross-correlation row vector between $\bm{z}(\bm{s}_{i})$ and $\bm{z}_{\mathcal{S}_{-i}}$, $\bm{H}_{\mathcal{S}_{-i}}$ is the $(n-1)\times(n-1)$ correlation matrix of $\bm{z}_{\mathcal{S}_{-i}}$, and the $j$-th element of $p$-dimensional vectors $\bm{\ell}$ and $\bm{u}$ are
@@
Scalable posterior computation under large spatial data
@@
$\mathcal{N}_{i}=\{\bm{s}_{i^{\prime}}|\mbox{$i^{\prime}$ is an index of $m$-nearest neighbor of $\bm{s}_{i}$}\}$, $\bm{H}_{\bm{s}_{i},\mathcal{N}_{i}}$ is the $n_{i}$-dimensional cross-correlation row vector between $\bm{z}(\bm{s}_{i})$ and $\bm{z}_{\mathcal{N}_{i}}$ with $n_{i}=|\mathcal{N}_{i}|(\leq m)$, and $\bm{H}_{\mathcal{N}_{i}}$ is the $n_{i}\times n_{i}$ correlation matrix of $\bm{z}_{\mathcal{N}_{i}}$. Under NNGP, which is a type of Vecchia approximation (Vecchia, 1988), the conditional distribution of $\bm{z}(\bm{s}_{i})$ can be expressed as
@@
In the calculation of $\tilde{\bm{\mu}}_{\bm{s}_{i}}$ and $\tilde{\bm{\Sigma}}_{\bm{s}_{i}}$ in the full conditional of $\bm{z}(\bm{s}_{i})$, $(n-1)$-dimensional matrix operations are required with the full Gaussian process, as seen in Section 3.2, but with the NNGP, we only need a $d_{i}$ ($<(n-1)$)-dimension matrix operations at most. Moreover, because only $m$ ($\ll n$)-dimensional matrix operations are required to sample $R$ and $\phi$ from their full conditional, computational cost can be considerably reduced compared to the full Gaussian process. Although the value of $m$ needs to be chosen based on the available computational resources, in many situations, relatively small values of $m$ (e.g., $m<100$) can achieve high accuracy. For a detailed discussion, see, for example, Katzfuss and Guinness (2021). In practice, in our method, even with a sample from a few hundreds to thousands of observations, the estimation accuracy remains almost the same, whereas the computation time is significantly reduced compared to the full Gaussian process.
Simulation study
Using a simulation study, this section demonstrates that the proposed approach successfully accounts for spatial correlation and correctly infers the dependence structure among spatially correlated multivariate mixed outcomes, including a comparison with Hoff (2007). To this end, we consider spatial Gaussian copula (3) with $n\in\{50,500,1000\}$ and $p\in\{6,9\}$, where $\phi\in\{0.05,0.25,0.5\}$ and $\bm{R}_{12}=0.5$, $\bm{R}_{14}=0.3$, $\bm{R}_{15}=0.2$, $\bm{R}_{23}=-0.2$, $\bm{R}_{24}=-0.3$, $\bm{R}_{35}=0.4$, $\bm{R}_{45}=-0.5$, and all others to zero for $j<j^{\prime}$. Location information $s_{i1}$ and $s_{i2}$ are generated from ${\rm Unif}(0,1)$. Using the spatial Gaussian copula, we generate a simulated dataset based on Smith (2021):
  1. Step 1.
    Calculate spatial correlation matrix $\bm{H}$ with location information $\bm{s}_{i}=(s_{i1},s_{i2})^{\top}$.
  2. Step 2.

### v3

# Semiparametric Copula Estimation for Spatially Correlated Multivariate Mixed Outcomes: Analyzing Visual Sightings of Fin Whales from a Line Transect Survey

@@ 1.1 Background and motivation
To model the interactions among variables, multivariate analysis is a ubiquitous tool, with many methodologies developed, such as factor analysis, principal component analysis (e.g., Manly and Alberto, 2016), graphical modeling (e.g., Jordan, 2004), and copulas (e.g., Joe, 2014). In particular, copulas are attractive because they decouple the dependence structure from the marginal distributions, allowing each component to be modeled separately (e.g., Joe, 2014). However, inference on copula parameters generally requires assumptions or estimates for the marginal distributions. To address this issue, Hoff (2007) proposed the extended rank likelihood, which enables inference on copula parameters without explicit parametric assumptions on the marginals and without requiring their estimation. This approach can be successfully applied even when outcomes are mixed (i.e., outcomes include both continuous and discrete variables). Following Hoff (2007), we refer to this framework as “semiparametric” because the dependence structure is modeled parametrically through the Gaussian copula, while the marginal distributions are left completely unspecified. Although these methods are useful for modeling the dependence structure, they assume that the multivariate observations are independent. However, beginning with the example of marine ecosystems mentioned above, in many applications such as epidemiology, climatology, medicine, and sociology, multivariate data with location information are often available, and spatial correlation must be appropriately considered; otherwise, the estimation of the dependence structure among outcomes could be severely biased.
@@
Related work in spatial modeling includes several studies on modeling multivariate spatial data. For example, Dey et al. (2022) and Krock et al. (2023) proposed efficient modeling (relatively high-dimensional) of multivariate spatial data for Gaussian outcomes, whereas Feng and Dean (2012) and Torabi (2014) considered multivariate models based on generalized linear mixed models. Furthermore, attempts have been made to use copulas in multivariate spatial modeling, such as Musafer et al. (2017), Krupskii et al. (2018), Krupskii and Genton (2019), and Gong and Huser (2022). Because the above methods consider the joint estimation of both marginal and dependence structures, one may lose the efficiency of the estimation by a large number of nuisance parameters in the marginal distributions, while the estimation of the dependence could be biased owing to potential misspecification of the marginal distributions. Hence, it is more reasonable to model the dependence structure directly, as in the proposed method.
@@
The remainder of this paper is organized as follows. Section 2 introduces the extended rank likelihood and semiparametric Bayesian inference for the dependence structure proposed by Hoff (2007). Section 3 proposes semiparametric Bayesian inference for the dependence structure in spatially correlated mixed outcomes and develops an efficient algorithm for computing the posterior using MCMC. The algorithm is then extended to a more scalable implementation using the nearest-neighbor Gaussian process. Section 4 presents the simulation results, including a comparison with the method of Hoff (2007), to validate the usefulness of the proposed method. Section 5 presents the application of the proposed method using the fin whale sighting data described in Section 1.2. Section 6 provides directions for future research. The R code for implementing the proposed method is available in the GitHub repository (https://github.com/t-momozaki/spBGC).
1.2 Motivating application: Fin whale sighting data
To motivate the proposed methodology, we describe a dataset for characterizing the Sourthern Ocean ecosystem. Antarctic krill (Euphausia superba) is a characteristic species of the Soruthern Ocean and exists within a narrow band of cold temperatures, as this species is known as cold adaptive species (Cui et al., 2025). This is a major prey item for a diverse suite of predators including whales, penguins, seals and fish and is an important fishery resource (Krafft et al., 2019). However, recent ocean warming has impacted population dynamics, with changes in sea ice dynamics-including a reduction-being observed (Kawaguchi et al., 2024).
The data was collected during the 2019 Area 48 Survey for Antarctic krill in the Scotia Sea of the Antarctic Peninsula (Krafft et al., 2021). During the krill survey period, visual sightings of fin whales (Balaenoptera physalus) were carried out along survey transects from three of the six participating vessels: R/V Kronprins Haakon (KPH), F/V Cabo de Hornos (CDH), and RRS Discovery (DIS), following the observation protocols detailed in Biuw et al. (2024). The observed data in the survey consisted of the number of fin whale sightings (groups and individuals) and krill acoustic data (recalculated to krill biomass; (details in the Supplementary Material of Krafft et al., 2021)) from vessels operating along the transects. The observation transects were split into 1-nm long segments, resulting in 3,833 segments based on 42 transects. For each segment, these data were summarized.
In addition, four environmental variables were included to investigate the association of environmental data with the biological community: sea surface temperature (SST, in Celsius), water depth (Depth, in meters), slope of the depth (Slope), and gradient of surface temperature (SST.grd). Water depth data were obtained from the ETOPO 1 bathymetric dataset (https://www.ncei.noaa.gov/products/etopo-global-relief-model), and SST data were obtained from the OISST dataset (https://www.ncei.noaa.gov/products/optimum-interpolation-sst), both extracted for the middle position of each 1-nm segment. The slope of the depth is the maximum rate of change in depth from that cell to its neighbors, calculated using the Slope tool of the Surface toolset in ArcGIS (https://desktop.arcgis.com/en/arcmap/10.3/tools/spatial-analyst-toolbox/an-overview-of-the-surface-tools.htm). The lower the slope value is, the flatter is the terrain, and vice versa. The gradient surface temperature (SST.grd) was calculated using the Slope tool applied to the SST.
Figure 2 shows the spatial plots of all six variables across the study region. The spatial plots reveal substantial spatial heterogeneity in both biological and environmental variables. To quantify the spatial correlation in the data, we compute Moran’s $I$ statistic (Moran, 1950), a widely used measure of spatial autocorrelation where values close to +1+1 indicate strong positive spatial correlation. Table 1 shows that all six variables exhibit positive spatial autocorrelation, with environmental variables (SST, Depth, SST.grd) showing particularly strong spatial correlation ($I>0.86$), while biological variables (Krill, Whale) exhibit moderate correlation ($I\approx 0.4$). This spatial correlation, combined with the mixed nature of the outcomes, makes conventional multivariate analysis methods inappropriate for inferring the dependence structure among these variables.
Table 1: Moran’s $I$ statistic for observed data. Values close to +1+1 indicate strong positive spatial autocorrelation.
Krill Whale SST Depth SST.grd Slope
$I_{\text{obs}}$ 0.410 0.442 0.999 0.989 0.981 0.868
Understanding the relationships between fin whales, their prey (krill), and environmental factors is essential for assessing the impacts of climate change on Antarctic marine ecosystems. However, this task presents several statistical challenges. First, the observed data consist of mixed outcomes: the number of fin whale sightings (count data) and krill biomass (continuous data). Second, the data exhibit strong spatial correlation due to the spatial distribution of both species and oceanographic features.
Our goal is to infer the dependence structure between krill biomass and fin whale sightings, as well as their relationships with environmental factors, while properly accounting for the spatial correlation in the data. This application demonstrates the need for the semiparametric spatial copula approach developed in this paper.
(a) Krill
(b) Whale
(c) SST
(d) Depth
(e) Slope
(f) SST.grd
Figure 2: Spatial distribution of the six variables in the fin whale sighting dataset: krill biomass (Krill), fin whale sightings (Whale), sea surface temperature (SST), water depth (Depth), slope of the depth (Slope), and gradient of surface temperature (SST.grd). Brighter/darker colors for SST and Depth indicate higher/lower temperatures and shallower/deeper water depths, respectively.
2 Extended rank likelihood for semiparametric copula estimation
@@
Suppose that we observe $p$-dimensional observations of mixed outcomes $\bm{y}_{i}=(y_{i1},\ldots,y_{ip})^{\top}$ for $i=1,2,\ldots,n$, where each variable could be a variety of outcomes, such as continuous, count, and ordered variables. Here, we are interested in the dependence structure of $p$ outcomes. The marginal distributions of each component in $\bm{y}_{i}$ are not of primary interest. Let $\bm{z}_{i}$ be a $p$-dimensional continuous latent variable for $\bm{y}_{i}$. Then, we consider the following Gaussian copula model:
@@
where $\bm{R}$ is the $p\times p$ correlation matrix and $\bm{I}_{n}$, $\otimes$, $F_{j}^{-1}$, and $\Phi(\cdot)$ denote the $n\times n$ identity matrix, Kronecker product, inverse of an unknown univariate cumulative distribution function, and cumulative distribution function of the standard normal distribution, respectively. The second line of (1) represents the relationship between the latent variable $z_{ij}$ and the observed outcome $y_{ij}$, following the notation of Hoff (2007).
Here, we focus on correlation matrix $\bm{R}$ as the dependence structure. Although the Gaussian copula model (1) separates the dependence structure $\bm{R}$ from the marginal distributions $F_{1},\ldots,F_{p}$, inference on $\bm{R}$ generally requires estimation or specification of these marginals. To overcome this difficulty, Hoff (2007) exploits the ordering constraints induced by the observed data. Since $F_{j}$ is a nondecreasing function, observing $y_{ij}<y_{i^{\prime}j}$ for any $i\neq i^{\prime}$ implies that $z_{ij}<z_{i^{\prime}j}$, regardless of the specific form of $F_{j}$. For discrete outcomes with tied observations (i.e., $y_{ij}=y_{i^{\prime}j}$), no ordering constraint is imposed between $z_{ij}$ and $z_{i^{\prime}j}$. Therefore, observing $\bm{y}=(\bm{y}_{1}^{\top},\bm{y}_{2}^{\top},\ldots,\bm{y}_{n}^{\top})^{\top}$ indicates that $\bm{z}$ must lie within the set
$D=\{\bm{z}\in\mathbb{R}^{pn}:\max(z_{kj}:y_{kj}<y_{ij})<z_{ij}<\min(z_{kj}:y_{ij}<y_{kj})\}.$
Since the set $D$ is determined solely by the ordering of the observed data, it does not depend on the specific form or continuity of the marginal distributions $F_{1},\ldots,F_{p}$. Based on this observation, Hoff (2007) defines the following extended rank likelihood:
$L(\bm{R})=\Pr(\bm{z}\in D|\bm{R})=\int_{D}\phi_{pn}(\bm{z};\bm{0},\bm{I}_{n}\otimes\bm{R})d\bm{z},$ (2)
where $\phi_{k}(\bm{x};\bm{\mu},\bm{\Sigma})$ is the $k$-dimensional multivariate normal density with a mean vector $\bm{\mu}$ and a variance-covariance matrix $\bm{\Sigma}$.
Note that the extended rank likelihood (2) depends only on the parameter of interest, $\bm{R}$, and not on the marginal distributions $\bm{F}=(F_{1},F_{2},\ldots,F_{p})^{\top}$. Although the marginal distributions implicitly influence the observed data $\bm{y}$ through the transformation in (1), the likelihood itself does not require their explicit specification or estimation. This also means that the discontinuity of $F_{j}$ for discrete outcomes does not affect inference on the dependence structure $\bm{R}$. Following Hoff (2007), we refer to this framework as “semiparametric” because the dependence structure is modeled parametrically through the Gaussian copula, while the marginal distributions are left completely unspecified.
The extended rank likelihood can be regarded as a type of marginal likelihood and can be derived in terms of the decomposition theorem on sufficient statistics. That is,
$p(\bm{y}|\bm{R},\bm{F})$ $=p(\bm{z}\in D,\bm{y}|\bm{R},\bm{F})$
@@
and to estimate $\bm{R}$, it is sufficient to use only $\Pr(\bm{z}\in D|\bm{R})$. The extended rank likelihood can also be seen as a multivariate version of the rank likelihood (Pettitt, 1982; Heller and Qin, 2001). For a more detailed discussion on the extended rank likelihood and sufficient statistics, see Hoff (2007), Section 5.
@@
3 Copula estimation under spatial correlation
3.1 Latent models with spatial correlation
Suppose that we obtain observations of mixed outcomes $\bm{y}_{\mathcal{S}}=(\bm{y}(\bm{s}_{1})^{\top},\bm{y}(\bm{s}_{2})^{\top},\ldots,\bm{y}(\bm{s}_{n})^{\top})^{\top}$ equipped with location information $\mathcal{S}=\{\bm{s}_{1},\bm{s}_{2},\ldots\bm{s}_{n}\}$. Here, $\bm{y}(\bm{s}_{i})=(y_{1}(\bm{s}_{i}),y_{2}(\bm{s}_{i}),\ldots,y_{p}(\bm{s}_{i}))^{\top}$ and $\bm{s}_{i}$ denotes a two-dimensional vector of longitude and latitude. These observations are characterized by spatial dependence and correlation, and nearby observations have similar properties. In such data, spatial correlations should be adequately considered; otherwise, the estimation of the dependence structure among outcomes would be severely biased. Therefore, we consider a hierarchical spatial model and latent multivariate Gaussian process. We define the following spatial Gaussian copula as a Gaussian copula model combined with a spatial Gaussian process to consider spatial correlation:
@@ 3.2 Posterior computation
[…]
@@
3.3 Scalable posterior computation under large spatial data
@@
[…]
Under NNGP, which is a type of Vecchia approximation (Vecchia, 1988), the conditional distribution of $\bm{z}(\bm{s}_{i})$ can be expressed as
@@
In the calculation of $\tilde{\bm{\mu}}_{\bm{s}_{i}}$ and $\tilde{\bm{\Sigma}}_{\bm{s}_{i}}$ in the full conditional of $\bm{z}(\bm{s}_{i})$, (n1)(n-1)-dimensional matrix operations are required with the full Gaussian process, as seen in Section 3.2, but with the NNGP, we only need a $d_{i}$ (<(n1)<(n-1))-dimension matrix operations at most. Moreover, because only $m$ ($\ll n$)-dimensional matrix operations are required to sample $R$ and $\phi$ from their full conditional, computational cost can be considerably reduced compared to the full Gaussian process. Specifically, the computational complexity of the NNGP is $O(nm^{2})$ per MCMC iteration, compared to $O(n^{3})$ for the full GP (Datta et al., 2016; Finley et al., 2019).
Although the value of $m$ needs to be chosen based on the available computational resources, in many situations, relatively small values of $m$ (e.g., $m<100$) can achieve high accuracy. For a detailed discussion, see, for example, Katzfuss and Guinness (2021). In practice, in our method, even with a sample from a few hundreds to thousands of observations, the estimation accuracy remains almost the same, whereas the computation time is significantly reduced compared to the full Gaussian process.
4 Simulation study
Using a simulation study, this section demonstrates that the proposed approach successfully accounts for spatial correlation and correctly infers the dependence structure among spatially correlated multivariate mixed outcomes, including a comparison with Hoff (2007). To this end, we consider spatial Gaussian copula (3) with $n\in\{50,500,1000\}$ and $p\in\{6,9\}$, where $\phi\in\{0.05,0.25,0.5\}$ and $\bm{R}_{12}=0.5$, $\bm{R}_{14}=0.3$, $\bm{R}_{15}=0.2$, $\bm{R}_{23}=-0.2$, $\bm{R}_{24}=-0.3$, $\bm{R}_{35}=0.4$, $\bm{R}_{45}=-0.5$, and all others to zero for $j<j^{\prime}$. Location information $\bm{s}_{i}=(s_{i1},s_{i2})^{\top}$ are generated independently, where both coordinates $s_{i1}$ and $s_{i2}$ are drawn from a uniform distribution on the unit square $0,1]\times[0,1]$, which represents a common scenario in spatial statistics where observations are randomly distributed across a study region. To examine the sensitivity of our results to the spatial configuration, we also conduct simulations using the actual spatial locations from the real data application; see Section A of the Supplementary Material.
Using the spatial Gaussian copula, we generate a simulated dataset based on Smith (2023):
  1. 1.
    Calculate spatial correlation matrix $\bm{H}$ with location information $\bm{s}_{i}=(s_{i1},s_{i2})^{\top}$, using the exponential correlation function for $\rho(\cdot;\phi)$ in (3).
  2. 2.